OpenStudy (anonymous):

Match each absolute value sentence with its equivalent compound sentence. Answer Potential Matches: A: |x - 2| = 6 1 : -6 < x - 2 < 6 B: |x - 2| < 6 2 : x - 2 > 6 or x - 2 < -6 C: |x - 2| > 6 3 : x - 2 = 6 or x - 2 = -6

OpenStudy (anonymous):

A: |x - 2| = 6 ... 3 : x - 2 = 6 or x - 2 = -6 B: |x - 2| < 6 ... 1 : -6 < x - 2 < 6 C: |x - 2| > 6 ... 2 : x - 2 > 6 or x - 2 < -6